OpenStudy (anonymous):

find the derivative y=ln((1+e^x)/(1-e^x)) plz help!!!

zepdrix (zepdrix):

\[\Large y=\ln\left(\frac{1+e^x}{1-e^x}\right)\]This thingy?

OpenStudy (anonymous):

yes @zepdrix

zepdrix (zepdrix):

Hmm so i guess applying a rule of logs before we differentiate might help:\[\Large \color{orangered}{\ln\left(\frac{a}{b}\right)\quad=\quad \ln(a)-\ln(b)}\]Understand how we can use this rule? :o

zepdrix (zepdrix):

We don't HAVE TO apply the rule if you don't want to :) It just allows us to avoid doing the quotient rule which can be a bit of a pain.

OpenStudy (anonymous):

it would be ln (1+e^x) - ln(1-e^x) right?

zepdrix (zepdrix):

yes, good.

zepdrix (zepdrix):

So ummm remember the rule for differentiating ln x? :)

OpenStudy (anonymous):

no, not really :(

zepdrix (zepdrix):

Hmm well here is the scoop :O\[\Large (\ln x)'\quad=\quad \frac{1}{x}\]Hopefully it looks familiar :D

OpenStudy (anonymous):

yes it does!! lol sorry xp

zepdrix (zepdrix):

When the contents of the log is `more than just x` we need to remember to apply the chain rule! \[\Large (\ln \color{royalblue}{stuff})'\quad=\quad \frac{1}{\color{royalblue}{stuff}}(\color{royalblue}{stuff})'\]

OpenStudy (anonymous):

ok so (1/1+e^x)(1e^x) yes??

zepdrix (zepdrix):

derivative of (1+e^x) gave you (e^x) when doing your chain? Ok looks good for the first log!

OpenStudy (anonymous):

yes I get e^x :o

zepdrix (zepdrix):

\[\Large y=\ln(1+e^x)-\ln(1-e^x)\] \[\Large \color{royalblue}{y'}\quad=\quad\frac{1}{1+e^x}(e^x)-\color{royalblue}{\left[\ln(1-e^x)\right]'}\]Ok so we took the derivative of that first log, how bout the other one? :O

OpenStudy (anonymous):

(-1/1-e^x)(-e^x) right?

OpenStudy (anonymous):

@zepdrix ^^

zepdrix (zepdrix):

Ah yes, good! :)

zepdrix (zepdrix):

You could probably simplify the expression by combining the fractions, but mehhh whatever.. depends whether or not your teacher wants you to do that.

OpenStudy (anonymous):

yey!!!!!!!!!! \(*_*)/

OpenStudy (anonymous):

and that is it!!!!???:O lol

zepdrix (zepdrix):

lol ya i guess so :)

OpenStudy (anonymous):

u guess?? lol :)

zepdrix (zepdrix):

Well there is simplification that `can` be done, but that's just the annoying stuff at the end :b As far as the differentiation is concerned, yes we're done! :D Just depends whether or not you want to simplify it.

OpenStudy (anonymous):

I simplified it:) lol and thank u so much!!!! I appreciate it:)

zepdrix (zepdrix):

yay \c:/

OpenStudy (anonymous):

lol u should be my actual tutor, u know everythin!